# Weighted Covariance Intersection for Range-based Distributed Cooperative Localization of Multi-Agent Systems

### v1

@@ IV-C WCI-based Correlated Update in DCL
In this subsection, continuing from the distributed cooperative localization framework introduced in Section III, we present the WCI-based correlated update process. Using agent $i$ as an example, assume it has obtained differential distances to the $N$ other agents and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{U}},\boldsymbol{R}_{ij}^{\mathcal{U}}\right\}$, $j=1,\ldots,N$, whose explicit expressions are given in (9) and (11), respectively. Since the noise of the differential distance measurements contains estimation errors of other agents’ states, which are correlated with the state of agent $i$, and these correlations are unknown, we can use CI, or its enhanced version WCI, introduced in Section III-B, to perform the correlated update process.
@@
Unlike EKF, where sequential and concurrent updates of multiple observations yield the same result, CI does not provide the same guarantee. In fact, CI ensures that the fusion of two data sources provides an optimal conservative upper bound, but it cannot guarantee the same for the fusion of multiple data sources [29]. Sequential updates of multiple observations lead to additional performance degradation and result in a more conservative upper bound. Therefore, for multiple distance measurements between agents, we combine them into a single observation vector for state updates.
$\boldsymbol{z}_{i}^{\mathcal{U}}=\left[\delta{d_{i1}^{\mathcal{U}}},\cdots,\delta{d_{iN}^{\mathcal{U}}}\right]^{\mathsf{T}}=\boldsymbol{H}_{i}^{\mathcal{U}}\delta\boldsymbol{x}^{i}+\boldsymbol{G}_{i}^{\mathcal{U}}\boldsymbol{v}_{i}^{\mathcal{U}}$ (21)
where the explicit expressions of $\boldsymbol{H}_{i}^{\mathcal{U}}$ and $\boldsymbol{G}_{i}^{\mathcal{U}}$ are given by:
$\boldsymbol{H}_{i}^{\mathcal{U}}$ $=[{\boldsymbol{H}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iN}^{\mathcal{U}}}^{\mathsf{T}}]^{\mathsf{T}}$ (22)
$\boldsymbol{G}_{i}^{\mathcal{U}}$ $=[{\boldsymbol{G}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{G}_{iN}^{\mathcal{U}}}^{\mathsf{T}}]^{\mathsf{T}}$
@@
$\boldsymbol{v}_{i}^{\mathcal{U}}$ $=\left[{\boldsymbol{v}_{i1}^{\mathcal{U}}}^{\mathsf{T}},\cdots,{\boldsymbol{v}_{iN}^{\mathcal{U}}}^{\mathsf{T}}\right]^{\mathsf{T}}$ (23)
$\boldsymbol{R}_{i}^{\mathcal{U}}$ $=\begin{bmatrix}\boldsymbol{R}_{i1}^{\mathcal{U}}&\cdots&\boldsymbol{C}_{i,1N}^{\mathcal{U}}\\ \vdots&\ddots&\vdots\\ {\boldsymbol{C}_{i,1N}^{\mathcal{U}}}^{\mathsf{T}}&\cdots&\boldsymbol{R}_{iN}^{\mathcal{U}}\end{bmatrix}$
Here, $\boldsymbol{C}_{i,1N}^{\mathcal{U}}$ denote the cross-covariance matrix between $\boldsymbol{v}_{i1}^{\mathcal{U}}$ and $\boldsymbol{v}_{iN}^{\mathcal{U}}$, which is generally non-zero since the states of agent 1 and agent $N$ are potentially correlated. However, $\boldsymbol{C}_{i,1N}^{\mathcal{U}}$ is not unavailable in practical distributed systems, as we do not track the correlations between agents. An alternative approach is to appropriately inflate the noise covariance matrix while ensuring that the nominal covariance matrix remains greater than the actual covariance matrix, thus maintaining consistency. We focus on the diagonal elements of $\boldsymbol{R}_{i}^{\mathcal{U}}$ and multiply the agent state error variance terms by 2 to obtain the nominal noise covariance matrix $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$, as follows:
$\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}=\begin{bmatrix}2\boldsymbol{P}_{p_{1}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{1}}^{d}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&2\boldsymbol{P}_{p_{N}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{N}}^{d}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}\end{bmatrix}$ (24)
@@
($j=1,\ldots,N$) are the diagonal matrices consisting of the error variances of agent $j$’s position and attitude. It is easy to observe that $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\succeq\boldsymbol{R}_{i}^{\mathcal{U}}$, meaning that the nominal covariance matrix $\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$ still provides a consistent estimate of the differential measurement vector. In addition, the nominal covariance matrix is free from the correlation items, thereby eliminating the need to track correlations and reducing the communication burden.
With the differential measurements between agents $\left\{\boldsymbol{z}_{i}^{\mathcal{U}},\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\right\}$, we can apply CI for state updates. A KF-style CI formulation for partial observations is provided in [16] as follows:
@@
$\boldsymbol{P}_{2}=\frac{1}{1-\omega}\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}$
$\boldsymbol{K}=\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{U}}}^{\mathsf{T}}\left({\boldsymbol{H}_{i}^{\mathcal{U}}}\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{U}}}^{\mathsf{T}}+\boldsymbol{P}_{2}\right)^{-1}$ (26)
$\delta{\boldsymbol{x}}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{U}}$
$\boldsymbol{P}_{i,k}=(\boldsymbol{I}-\boldsymbol{K}{\boldsymbol{H}_{i}^{\mathcal{U}}})\boldsymbol{P}_{1}$
Here, $\omega$ is the weight parameter, which is determined by a specific optimization criterion in (16). We find that both the classical criteria of minimizing the trace or the determinant of the fused covariance matrix perform poorly in this problem. Instead, we choose WCI as presented in (20), with the primary challenge lying in the selection of the weighting matrix.
@@
In the complete state estimation procedure for agents, inertial navigation serves as the primary method for state propagation, while distance measurements are used as a supplementary tool for state updates. If the error in inertial navigation can be minimized as much as possible, the overall state estimation error in cooperative localization will decrease, thereby enhancing the agent’s ability to navigate autonomously in scenarios with insufficient distance measurements. In summary, the optimization goal in the state update phase can be framed as minimizing the inertial navigation error.
We consider the propagation law of inertial navigation errors. Generally, agents are equipped with low cost micro-electromechanical system (MEMS) sensors, so we do not account for factors such as the earth’s rotation, the variation of frame $\mathcal{N}$, or changes in gravity. As a result, a simplified error state model can be expressed as [24]:
@@
We make a simple assumption that inertial navigation position error accumulates similarly to the case where agent $i$ moves with constant velocity in a straight line over a time period of $T_{a}$. The inertial navigation position error at time $T_{a}$ can then be approximately written as:
@@
where the explicit expression for $\boldsymbol{S}$ is given by (30). We aim to minimize the mean square error of $\prescript{I}{}{\delta{\boldsymbol{p}}}_{i,T_{a}}^{n}$, which corresponds to minimizing the trace of the product of $\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ and initial error-state covariance matrix $\boldsymbol{P}_{i,0}$ as illustrated in (31).
@@
where $\boldsymbol{W}_{i,k}$ is determined by (30) and (32) with predicted state estimates and IMU outputs at time step $k$. The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can be used to correct nominal state estimates as well. The entire procedures of the WCI-based correlated update are summarized in Algorithm 1.
Algorithm 1 WCI-based Correlated Update
1:
2:
@@
  • Distance measurements with other agents $\{d_{ij}^{\mathcal{U}}\}$
  • Other agents’ position and attitude estimates $\prescript{\mathcal{I}}{}{\hat{\boldsymbol{p}}_{j,k|k-1}^{n}}$ and $\hat{\boldsymbol{q}}_{j,k|k-1}^{n}$, along with their nominal variances $\{\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2}\sigma_{p_{j,z}}^{2}\}$ and $\{\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2}\sigma_{\phi_{j,z}}^{2}\}$.
3:State estimate after update $\{\hat{\boldsymbol{x}}_{i,k},\boldsymbol{P}_{i,k}\}$
4:Calculate differential measurements with the inflated nominal covariance matrices $\{\boldsymbol{z}_{i}^{\mathcal{U}},\boldsymbol{R}_{i}^{\mathcal{U}^{\prime}}\}$ according to (17), (21), (24) and (25).
5:Determine the weighting matrix $\boldsymbol{W}_{i,k}$ by (30) and (32), replacing ${\boldsymbol{f}}_{k,i}^{b}$ with $\tilde{\boldsymbol{f}}_{k,i}^{b}$.
6:Execute data fusion using WCI by (26) and (33).
7:Perform error feedback to correct the predicted state estimate.
Furthermore, we have described the independent and correlated update procedures for distance measurements between agents and anchors, and between agents, using EKF and WCI, respectively. When both types of measurements are available, they can be processed sequentially. The updated states and covariance matrix are then prepared for the next state prediction procedure.
@@ V Simulation Results
In this section, we conduct simulation experiments to validate the superiority of WCI over classical CI in cooperative localization. We also compare the proposed DCL framework with the CCL framework to demonstrate its superiority in terms of robustness, as well as reduced computation and communication burdens.
Consider a 3D cooperative localization system where all agents initially remain static, with random positions and random attitudes (with roll and pitch set to zero). They then perform the following motion:
@@
The total motion lasts 120 seconds, encompassing a variety of typical movements and sufficiently activating the observability of all state components. To prevent the agents’ states from diverging, we set up four anchors with known positions, located at $100,0]^{\mathsf{T}}\text{m}, [100,100,$$100]^{\mathsf{T}}\text{m}, [100,100,$$100,0]^{\mathsf{T}}\text{m}$ and $100,-100,100]^{\mathsf{T}}\text{m}$, respectively. The IMU parameters are set according to the ADIS16465 [30], which is a typical precise MEMS IMU. Regarding ranging techniques, we choose UWB, which can achieve centimeter-level ranging accuracy under line-of-sight conditions [31]. Unless otherwise specified, the remaining simulation parameters are set as shown in Table II. The trajectories of the agents and the positions of the anchors are depicted in Fig. 4.
We assume that agent 8 is connected to all four anchors, while agents 2-7 are connected to only two of the four anchors, which is common in practice due to communication range limitations or obstructions. However, agent 1 cannot connect to any anchors and can only correct its state by cooperating with other agents.
@@
$\sigma_{p_{0}}$ $0.3\,\text{m}$ Standard deviation of each agent’s initial position
$\sigma_{v_{0}}$ $0.1\,\text{m/s}$ Standard deviation of each agent’s initial velocity
$\sigma_{\phi_{0}}$ $3^{\circ}$ Standard deviation of each agent’s initial attitude
@@
Figure 4: Anchor distributions and trajectories of the agents in the simulation experiments
@@
    NCL: No cooperative localization, where each agent corrects its state using only distance measurements from anchors. This method serves as a baseline to highlight the impact of inter-agent measurements and cooperation.
@@
    EKF: A CCL method where the computing center [^2] Unless otherwise specified, we assume that all IMU and distance measurements can be transmitted to the computing center without packet loss. maintains a large covariance matrix for all agents’ states and utilizes EKF for updates using all available measurements. Since it tracks the correlations between agents, it is considered the performance ceiling for cooperative localization.
@@
    CI-trace: A DCL method where each agent runs its own filter and uses trace-based CI for updates with measurements between agents.
@@
    CI-det: A DCL method where each agent runs its own filter and uses determinant-based CI for updates with measurements between agents.
We evaluate the state estimation results based on the state components of position, velocity, attitude, gyroscope bias, and accelerometer bias, respectively. The evaluation metrics used include root mean square error (RMSE), standard deviation (STD), and normalized estimation error squared (NEES). RMSE reflects the error between the estimated states and the true values, while STD indicates how well the nominal covariance matrix represents the uncertainty. NEES measures the relative magnitude of the estimation error compared to its covariance matrix, reflecting the consistency of the estimate [^3] Ideally, for a random vector of dimension $n$, the theoretical value of NEES is $n$; if NEES is significantly greater or smaller than $n$, the estimator is either overly pessimistic or optimistic. Let $\boldsymbol{y}$ denote the state component under consideration. At the time step $k$, the state estimate and true value are denoted as $\hat{\boldsymbol{y}}_{k}$ and $\boldsymbol{y}_{k}$, respectively, with the nominal covariance matrix given by $\boldsymbol{P}_{y,k}$. The formulas for these three metrics are as follows:
@@
$\operatorname{STD}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}\operatorname{tr}(\boldsymbol{P}_{y,k})}$ (35)

### v2

@@ IV-B Concurrent Update Strategy
From this subsection, continuing from the distributed cooperative localization framework introduced in Section III, we present the WCI-based correlated update process. Using vehicle $i$ as an example, assume it has obtained differential distances to the $N$ other vehicles and their corresponding covariance matrices be denoted as $\left\{\delta d_{ij}^{\mathcal{V}},\boldsymbol{R}_{ij}^{\mathcal{V}}\right\}$, $j=1,\ldots,N$, whose explicit expressions are given in (9) and (11), respectively. Since the noise of the differential distance measurements contains estimation errors of other vehicles’ states, which are correlated with the state of vehicle $i$ and these correlations are unknown, we can use CI, or its enhanced version WCI, introduced in Section IV-A, to perform the correlated update process.
@@
Unlike EKF, where sequential and concurrent updates of multiple observations yield the same result, CI does not provide the same guarantee. In fact, CI ensures that the fusion of two data sources provides an optimal conservative upper bound, but it cannot guarantee the same for the fusion of multiple data sources [32]. Sequential updates of multiple observations lead to additional performance degradation and result in a more conservative upper bound. Therefore, for multiple distance measurements between vehicles, we combine them into a single observation vector for state updates.
$\boldsymbol{z}_{i}^{\mathcal{V}}=\left[\delta{d_{i1}^{\mathcal{V}}},\cdots,\delta{d_{iN}^{\mathcal{V}}}\right]^{\mathsf{T}}=\boldsymbol{H}_{i}^{\mathcal{V}}\delta\boldsymbol{x}^{i}+\boldsymbol{G}_{i}^{\mathcal{V}}\boldsymbol{v}_{i}^{\mathcal{V}}$ (20)
where the explicit expressions of $\boldsymbol{H}_{i}^{\mathcal{V}}$ and $\boldsymbol{G}_{i}^{\mathcal{V}}$ are given by:
$\boldsymbol{H}_{i}^{\mathcal{V}}$ $=[{\boldsymbol{H}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{H}_{iN}^{\mathcal{V}}}^{\mathsf{T}}]^{\mathsf{T}}$ (21)
$\boldsymbol{G}_{i}^{\mathcal{V}}$ $=[{\boldsymbol{G}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{G}_{iN}^{\mathcal{V}}}^{\mathsf{T}}]^{\mathsf{T}}$
@@
$\boldsymbol{v}_{i}^{\mathcal{V}}$ $=\left[{\boldsymbol{v}_{i1}^{\mathcal{V}}}^{\mathsf{T}},\cdots,{\boldsymbol{v}_{iN}^{\mathcal{V}}}^{\mathsf{T}}\right]^{\mathsf{T}}$ (22)
$\boldsymbol{R}_{i}^{\mathcal{V}}$ $=\begin{bmatrix}\boldsymbol{R}_{i1}^{\mathcal{V}}&\cdots&\boldsymbol{C}_{i,1N}^{\mathcal{V}}\\ \vdots&\ddots&\vdots\\ {\boldsymbol{C}_{i,1N}^{\mathcal{V}}}^{\mathsf{T}}&\cdots&\boldsymbol{R}_{iN}^{\mathcal{V}}\end{bmatrix}$
Here, $\boldsymbol{C}_{i,1N}^{\mathcal{V}}$ denote the cross-covariance matrix between $\boldsymbol{v}_{i1}^{\mathcal{V}}$ and $\boldsymbol{v}_{iN}^{\mathcal{V}}$, which is generally non-zero since the states of vehicle 1 and vehicle $N$ are potentially correlated. However, $\boldsymbol{C}_{i,1N}^{\mathcal{V}}$ is not unavailable in practical distributed systems, as we do not track the correlations between vehicles. An alternative approach is to appropriately inflate the noise covariance matrix while ensuring that the nominal covariance matrix remains greater than the actual covariance matrix, thus maintaining consistency. We focus on the diagonal elements of $\boldsymbol{R}_{i}^{\mathcal{V}}$ and multiply the vehicle state error variance terms by 2 to obtain the nominal noise covariance matrix $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$, as follows:
$\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}=\begin{bmatrix}2\boldsymbol{P}_{p_{1}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{1}}^{d}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&2\boldsymbol{P}_{p_{N}}^{d}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 1}&\cdots&\boldsymbol{0}_{3\times 3}&2\boldsymbol{P}_{\phi_{N}}^{d}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0&\cdots&\boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&\sigma_{r}^{2}\end{bmatrix}$ (23)
@@
($j=1,\ldots,N$) are the diagonal matrices consisting of the error variances of vehicle $j$’s position and attitude. It is easy to observe that $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\succeq\boldsymbol{R}_{i}^{\mathcal{V}}$, meaning that the nominal covariance matrix $\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$ still provides a consistent estimate of the differential measurement vector. In addition, the nominal covariance matrix is free from the correlation items, thereby eliminating the need to track correlations and reducing the communication burden.
With the differential measurements between vehicles $\left\{\boldsymbol{z}_{i}^{\mathcal{V}},\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\right\}$, we can apply CI for state updates. A KF-style CI formulation for partial observations is provided in [20] as follows:
@@
$\boldsymbol{P}_{2}=\frac{1}{1-\omega}\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}$
$\boldsymbol{K}=\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{V}}}^{\mathsf{T}}\left({\boldsymbol{H}_{i}^{\mathcal{V}}}\boldsymbol{P}_{1}{\boldsymbol{H}_{i}^{\mathcal{V}}}^{\mathsf{T}}+\boldsymbol{P}_{2}\right)^{-1}$ (25)
$\delta{\boldsymbol{x}}_{i,k}=\boldsymbol{K}\boldsymbol{z}_{i}^{\mathcal{V}}$
$\boldsymbol{P}_{i,k}=(\boldsymbol{I}-\boldsymbol{K}{\boldsymbol{H}_{i}^{\mathcal{V}}})\boldsymbol{P}_{1}$
Here, $\omega$ is the weight parameter, which is determined by a specific optimization criterion in (16).
In the context of our problem, the state vector includes components with large scale and high observability, such as position, as well as components with small scale and low observability, such as attitude and IMU biases. Traditional criteria, such as CI-trace, typically reduce the uncertainty of large scale and highly observable components like position, but at the cost of significantly increasing the uncertainty of other state components. This imbalance ultimately degrades the overall estimation performance. In contrast, CI-det focuses more on optimizing the smaller-scale variables, giving more attention to those components with low scale. As a result, both the traditional criterion of minimizing the trace or the determinant of the fused covariance matrix perform poorly in the 3D cooperative localization task. Instead, we choose WCI as presented in (19), with the primary challenge lying in the selection of the weighting matrix.
IV-C Weighting Matrix Selection
@@
Inertial navigation serves as the primary method for state propagation in the vehicle state estimation procedure, while distance measurements are used as a supplementary tool for state correction. If the error in inertial navigation can be minimized as much as possible, the overall state estimation error in cooperative localization will decrease, thereby enhancing the vehicle’s autonomous navigation ability even in scenarios without sufficient distance measurements. In summary, the optimization goal in the state update phase can be framed as minimizing the inertial navigation error. Additionally, inertial navigation error is related to all state components, and therefore, optimizing these errors can provide guidance for achieving a balanced ratio between different state components.
We consider the propagation law of inertial navigation errors. Generally, vehicles are equipped with low cost micro-electromechanical system (MEMS) sensors, so we do not account for factors such as the earth’s rotation, the variation of frame $\mathcal{N}$, or changes in gravity. As a result, a simplified error state model can be expressed as [28]:
@@
We make a simple assumption that inertial navigation position error accumulates similarly to the case where vehicle $i$ moves with constant velocity in a straight line over a time period of $T_{a}$. The inertial navigation position error at time $T_{a}$ can then be approximately written as:
@@
where the explicit expression for $\boldsymbol{S}$ is given by (30). We aim to minimize the mean squared error of $\prescript{I}{}{{\boldsymbol{p}}}_{i,T_{a}}^{n}$, which corresponds to minimizing the trace of the product of $\boldsymbol{S}^{\mathsf{T}}\boldsymbol{S}$ and initial error-state covariance matrix $\boldsymbol{P}_{i,0}$ as illustrated in (31).
@@
where $\boldsymbol{W}_{i,k}$ is determined by (30) and (32) with predicted state estimates and IMU outputs at time step $k$. The estimated error-state $\delta\boldsymbol{x}_{i,k}$ can be used to correct nominal state estimates as well. The entire procedures of the WCI-based correlated update are summarized in Algorithm.
Algorithm WCI-based Correlated Update Method
@@
  • Distance measurements with other vehicles $\{d_{ij}^{\mathcal{V}}\}$
  • Other vehicles’ position and attitude estimates $\prescript{\mathcal{I}}{}{\hat{\boldsymbol{p}}_{j,k|k-1}^{n}}$ and $\hat{\boldsymbol{q}}_{j,k|k-1}^{n}$, along with their nominal variances $\{\sigma_{p_{j,x}}^{2},\sigma_{p_{j,y}}^{2}\sigma_{p_{j,z}}^{2}\}$ and $\{\sigma_{\phi_{j,x}}^{2},\sigma_{\phi_{j,y}}^{2}\sigma_{\phi_{j,z}}^{2}\}$.
State estimate after update $\{\hat{\boldsymbol{x}}_{i,k},\boldsymbol{P}_{i,k}\}$
Calculate differential measurements with the inflated nominal covariance matrices $\{\boldsymbol{z}_{i}^{\mathcal{V}},\boldsymbol{R}_{i}^{\mathcal{V}^{\prime}}\}$ according to (17), (20), (23) and (24).
Determine the weighting matrix $\boldsymbol{W}_{i,k}$ by (30) and (32), replacing ${\boldsymbol{f}}_{k,i}^{b}$ with $\tilde{\boldsymbol{f}}_{k,i}^{b}$.
Execute data fusion using WCI by (25) and (33).
Perform error feedback to correct the predicted state estimate.
Furthermore, we have described the independent and correlated update procedures for distance measurements between vehicles and anchors, and between vehicles, using EKF and WCI, respectively. When both types of measurements are available, they can be processed sequentially. The updated states and covariance matrix are then prepared for the next state prediction procedure.
@@ V Simulation Results
In this section, we conduct simulation experiments to validate the superiority of WCI over traditional CI in 3D cooperative localization. We also compare the proposed DCL framework with the CCL framework to demonstrate its superiority in terms of robustness, as well as reduced computation and communication burdens.
Consider a 3D cooperative localization system where all vehicles initially remain static, with random positions and random attitudes (with roll and pitch set to zero). They then perform the following motion:
@@
The total motion lasts 120 seconds, encompassing a variety of typical movements and sufficiently activating the observability of all state components. To prevent the vehicles’ states from diverging, we set up four anchors with known positions, located at $100,0]^{\mathsf{T}}\text{m}, [100,100,$$100]^{\mathsf{T}}\text{m}, [100,100,$$100,0]^{\mathsf{T}}\text{m}$ and $100,-100,100]^{\mathsf{T}}\text{m}$, respectively. The IMU parameters are set according to the ADIS16465 [33], which is a typical precise MEMS IMU. Regarding ranging techniques, we choose UWB, which can achieve centimeter-level ranging accuracy under line-of-sight conditions [34]. Unless otherwise specified, the remaining simulation parameters are set as shown in Table II. The trajectories of the vehicles and the positions of the anchors are depicted in Fig. 5.
We assume that vehicle 8 is connected to all four anchors, while vehicles 2-7 are connected to only two of the four anchors, which is common in practice due to communication range limitations or obstructions. However, vehicle 1 cannot connect to any anchors and can only correct its state by cooperating with other vehicles.
@@
$\sigma_{p_{0}}$ $0.3\,\text{m}$ Standard deviation of each vehicle’s initial position
$\sigma_{v_{0}}$ $0.1\,\text{m/s}$ Standard deviation of each vehicle’s initial velocity
$\sigma_{\phi_{0}}$ $3^{\circ}$ Standard deviation of each vehicle’s initial attitude
@@
Figure 5: Anchor distributions and trajectories of the vehicles in the simulation experiments
@@
    NCL: No cooperative localization, where each vehicle corrects its state using only distance measurements from anchors. This method serves as a baseline to highlight the impact of inter-vehicle measurements and cooperation.
@@
    CCL EKF: A CCL method where the computing center [^2] Unless otherwise specified, we assume that all IMU and distance measurements can be transmitted to the computing center without packet loss. maintains a large covariance matrix for all vehicles’ states and utilizes EKF for updates using all available measurements. Since it tracks the correlations between vehicles, it is considered the performance ceiling for cooperative localization.
@@
    DCL CI-trace: A DCL method where each vehicle runs its own filter and uses trace-based CI for updates with measurements between vehicles.
@@
    DCL CI-det: A DCL method where each vehicle runs its own filter and uses determinant-based CI for updates with measurements between vehicles.
We evaluate the state estimation results based on the state components of position, velocity, attitude, gyroscope bias, and accelerometer bias, respectively. The evaluation metrics used include root-mean-squared-error (RMSE), root-mean-trace-error (RMTE) [23], and normalized estimation error squared (NEES). RMSE reflects the error between the estimated states and the true values, while RMTE indicates how well the nominal covariance matrix represents the uncertainty. NEES measures the relative magnitude of the estimation error compared to its covariance matrix, reflecting the consistency of the estimate [^3] Ideally, for a random vector of dimension $n$, the theoretical value of NEES is $n$; if NEES is significantly greater or smaller than $n$, the estimator is either overly pessimistic or optimistic. Let $\boldsymbol{y}$ denote the state component under consideration. At the time step $k$, the state estimate and true value are denoted as $\hat{\boldsymbol{y}}_{k}$ and $\boldsymbol{y}_{k}$, respectively, with the nominal covariance matrix given by $\boldsymbol{P}_{y,k}$. The formulas for these three metrics are as follows:
@@
$\operatorname{RMTE}_{y,k}=\sqrt{\frac{1}{N_{\text{sim}}}\sum_{n=1}^{N_{\text{sim}}}\operatorname{tr}(\boldsymbol{P}_{y,k})}$ (35)